OpenStudy (yanni):

5(v-2)=5v+5-2(-4v-4)

OpenStudy (lgbasallote):

distribute... 5(v) - 5(2) = 5v + 5 -2(-4v) - 2(-4) need more hints? or are you good?

OpenStudy (anonymous):

Did you get it? or you need more help

OpenStudy (yanni):

im on a different question now. but thx

OpenStudy (yanni):

its -9/4

OpenStudy (lgbasallote):

i dont think it should come out negative...nor 9/4....

OpenStudy (lgbasallote):

5v - 10 = 5v + 5 + 8v + 8 5v -5v - 8v = 10 + 5 + 8 -8v = 23 v = -23/8 well it is negative...but not 9/4 lol